LJour2: reference card version 1.0

This is a list of commands provided by the LaTEX style file LJour2. For a complete description of these commands please refer to the user’s guide.

Using the style file

tex &cljour2 <filename> or tex &pljour2 <filename>

For Unix installations: remember to quote or to escape the &-chracter as it has a special meaning for the shell. Either use tex '&cljour2' <filename> or

tex \&pljour2 <filename>

Documentstyle

cljour2 or pljour2

Styleoptions

referee  footinfo  leqno  bibay  deutsch  german  francais  amssymb

The beginning of an article

\documentstyle[<style options>]{<docstyle>}
<macro definitions>
\journalname{<name of journal>}
\renewcommand{\abstractname}{<heading for abstract>.}
\renewcommand{\subclassname}{{\it <heading>.}}
\begin{document}
\headnote{<text>}
\title{<title of article>}
\titlerunning{<short form of title>}
\subtitle{<subtitle of article>}
\author{<name of first author>\inst{<number>}
\and <name of second author>\inst{<number>} ...}
\authorrunning{<name of first author> et al.}
\institute{<first address> \and <second address>}
\thanks{<text of footnote>}
\thanks{{\it Present address:\/} <address>}
\mail{<name>}
\date{Received <date>; accepted <date>}
\dedication{<dedication>}
\maketitle
\begin{abstract}  <abstract>  \end{abstract}
\keywords{<keyword -- keyword -- ...>}
\subclass{<classification>}
\transtitle{<title in English>}
\begin{transsummary} <English translation of summary>
\end{transsummary}
\transkey  {<key words in English>}
\strich

Footnote mark separator

\fnmsep

Headings

\section{<title of section>}
\subsection{<title of subsection>}
\subsubsection{<title of subsection>}
\paragraph{<title of paragraph>}

Special typefaces

\vec{<letter>}
\tens{<letter>}
{\Bbb <symbol>}
{\sc <text>}

Footnotes

\footnote{<text>}

Lists

\begin{description}[<largelabel>]
\item[<label>] <text>
\end{description}

Paragraphs in small print

\begin{petit}   <text>   \end{petit}

Mini-environments

\begin{<name of environment>}[<optional material>]
<text for that environment>
\end{<name of environment>}

Defining new environments

\newtheorem{<env_nam>}{<caption>}[<within>]
{<cap_font>}{<body_font>}

or

\newtheorem{<env_nam>}[<num_like>]{<caption>}
{<cap_font>}{<body_font>}

Listings of computer programs

\verb|<code>|
\begin{verbatim}
<code>
\end{verbatim}

Figures

The ordinary form of the environment is for one column figures and the *-form produces double column ones.

One figure

\begin{figure}
\picplace{<height> cm}
\caption[]{<text>}
\end{figure}

Two figures side by side

\begin{figure}
\picplace{<height>cm}
\firstcaption{<legend of first figure>}
\secondcaption{<legend of second figure>}
\end{figure}

Two figures side by side with legend one below the other

\begin{figure}
\picplace{<height> cm}
\caption{<legend of first figure>}
\caption{<legend of second figure>}
\end{figure}

Double column figure with legend attached at its side

\begin{figure*}
\sidecaption
\mpicplace{<width> cm}{<height> cm}
\caption{<legend of figure>}
\end{figure*}

Tables

The ordinary form of the environment is for one column tables and the *-form produces double column ones.

Coded with LaTEX

\begin{table}
\caption[ ]{<table caption>}

\begin{tabular}{<alignment specs>}
<input for table>
\end{tabular}
\end{table}

Not coded with LaTEX

\begin{table}
\caption[]{<text of caption>}
\picplace{<height> cm}
\end{table}

Acknowledgements

\begin{acknowledgement} <text> \end{acknowledgement}

Appendices

\appendix

References

\cite{<key>}

By number or by letter and number

\begin{thebibliography}[<label>]{<title>}
<bibliographic entries>
\end{thebibliography}

For details please see LaTEX manual.

Author-year system

\begin{thebibliography}[<label>]{<title>}
\bibitem{}{<key>}{<citation>} <bibliographic entry>
...
\end{thebibliography}

Biography

\begin{biography}{<author's name>}
<text of biography>
\end{biography}

Springer International